The dress was marked down 20% of its original price for the sale, and it still sold for $120. What would it sell for if the markdown of the original price was only 10%?
Answer by ikleyn(52790)

.
The dress was marked down 20% of its original price for the sale,
and it still sold for $120. What would it sell for if the markdown
of the original price was only 10%?
~~~~~~~~~~~~~~~~~

Let x be the original price (before the 20% markwown).

The problem says

    (1-0.2)x = 120  dollars,

or  0.8x = 120;  x = 120/0.8 = 150 dollars.


Thus the original price of the dress was $150.


Hence, at 10% markdown, the celling price would be 150 - 0.1*150 = 150 - 15 = 135 dollars.    ANSWER

Solved.